AutoSave OFF Ô sv5- Group Project 1 Part B-- Spring 2020_for distribution Home Insert Draw Page Layout Formulas Data Review VAutoSave OFF A Svor Group Project 1 Part B-- Spring 2020_for distribution Home Insert Draw Page Layout Formulas Data Review VI need help making a general journal from those transactions. I also need help with the closing entries at the bottom of the journal2 Group Project 1 Part B-- Spring 2020_for distribution-Saved AutoSave On Home Insert svo Page Layout Draw Formulas Data Revi
Answer #1

Journal

S.No. Date Account Titles Debit Credit
1 Jan-01 Patent/ Intangible asset $         32,000
     Cash $         32,000
(Purchased patent for cash)
2 Mar-15 Long-term Assets $         18,500
     Cash $         18,500
(Purchased long term assets)
3 Apr-01 Prepaid Rent $         36,000
      Cash $         36,000
(Paid rent for 12 months)
4 Apr-01 Cash $         60,000
      Common Stock $         60,000
(Issued 1000 common shares)
5a Apr-01 Accounts Receivable $       120,000
       Sales Revenue $       120,000
(Sale of inventory on account)
5b Apr-01 Cost of Goods Sold $         16,000
      Inventory $         16,000
(Cost of goods sold taken out of inventory)
5c Apr-01 Sales Commission $         10,000
       Cash $         10,000
(Paid sales commission)
6 Apr-05 Research and Development Expense $         13,400
       Cash $         13,400
(Paid Research and Development costs)
7 May-01 Cash $         24,000
      Unearned Revenue $         24,000
(Advance received for services)
8 May-01 Cash $         65,000
      Accounts Receivable $         65,000
(Collected Accounts receivable)
9 Jun-30 Income Tax Payable $            4,300
       Cash $            4,300
(Paid last year income tax liability)
10 Jul-31 Inventory $         15,000
      Cash $         12,000
      Accounts Payable $            3,000
(Purchased inventory)
11 Jul-31 Cash $         20,000
      Accounts Receivable $         20,000
(Collected Accounts receivable)
12a Sep-30 Accounts Receivable $            7,000
      Cash $            7,000
(Purchased Supplies)
12b Oct-01 Supplies $            7,000
      Accounts Receivable     $            7,000
(Rectification of error)
13 Oct-01 Marketable Securities $            6,000
      Cash $            6,000
(Purchased marketable securities for cash)
14 Oct-01 Cash $         20,000
      Notes Payable $         20,000
(Borrowed from local bank)
15 Dec-01 Salaries Expense $         34,000
      Cash $            4,000
      Salaries Payable $         30,000
(Salaries expense)
16 Dec-31 Income Tax Expense $            5,500
       Cash $            3,500
       Income tax Payable $            2,000
(Income tax expense for the year)
17 Dec-31 Rent Expense $         27,000 =36000*9/12
      Prepaid Rent $         27,000
(Recorded rent expense for 9 months)
18 Dec-31 Unearned Revenue $            4,000 =24000/6*1
      Service Revenue $            4,000
(Recorded revenue for 1 month)
19 Dec-31 Supplies Expense $            5,000 =7000-2000
      Supplies $            5,000
(Supplies consumed during the year)
20 Dec-31 Interest Expense $               500 =20000*10%*3/12
      Interest Payable $               500
(Recorded interest expense for 3 months)

Closing Entries

S.No. Date Account Titles Debit Credit
1 Dec-31 Sales Revenue $       120,000
Service Revenue $            4,000
     Income Summary $       124,000
(To close revenue accounts)
2 Dec-31 Income Summary $       111,400
Cost of Goods Sold $         16,000
Sales Commission $         10,000
Research & Development Costs $         13,400
Salaries Expense $         34,000
Supplies Expense $            5,000
Rent Expense $         27,000
Interest Expense $               500
Income Tax $            5,500
(To close expense accounts)
3 Dec-31 Income Summary $         12,600
     Retained Earnings $         12,600
(To close income summary account)